OpenStudy (anonymous):

Which function is the inverse of f(x) = 2x + 3?

OpenStudy (anonymous):

inverse of f(x)= x/2-3/2 you solve the equation for x to find what will make it 1

jimthompson5910 (jim_thompson5910):

First replace f(x) with y. Then swap x and y. Finally, solve for y to get your inverse function. \[\Large f(x) = 2x+3\] \[\Large y = 2x+3\] \[\Large x = 2y+3\] I'll let you finish up